OpenStudy (anonymous):

I don't know where to begin. Given f(x) = (2x - 2) / (x^2 + x - 2) and g(x) = (a) / (b + x) such that g(x) = f(x) wherever f is defined, find the values of a and b.

OpenStudy (amistre64):

factor ....

OpenStudy (anonymous):

I end up with \[f(x) = \frac{ 2 }{ 2+x } = g(x) = \frac{ a }{ b+x }\] but a and b can't be the same values, can they?

OpenStudy (amistre64):

there is nothing to say that cant be the same value

OpenStudy (anonymous):

If they were the same value, I think my teacher would have written g(x) = (a)/(a+x)

OpenStudy (amistre64):

a is some value, b is some value. since some value can be the same as some value .... then its possible for a to equal b, unless stated that a not equal b .... which is not states

OpenStudy (anonymous):

Ok that makes sense. My teacher has been known to trick us. Thanks.

OpenStudy (amistre64):

youre welcome